OpenStudy (highschoolmom2010):

The diameter of the cylinder is 1 yd. The height is 4 yd.

OpenStudy (anonymous):

What do you want to know? The cylinder's volume?

OpenStudy (anonymous):

@highschoolmom2010

OpenStudy (highschoolmom2010):

yes i was about to post that

OpenStudy (anonymous):

The formula to find a cylinder's volue is V=π^2*d*h. So the answer is: π^2*1*4 =4π^2

OpenStudy (anonymous):

@highschoolmom2010

OpenStudy (highschoolmom2010):

but if the diameter is 1 yard dont i have to split 1 yard in half to get the radius??

OpenStudy (highschoolmom2010):

my book gave me the formula : \[V=\pi r^2h\]

OpenStudy (nirmalnema):

h=4yard r=0.5yard \[\pi R ^{2} h\] 22/7*0.5*0.5.4 0.25*4*22/7 =1*22/7 =22/7 =pi

OpenStudy (highschoolmom2010):

thats what i thought i had to do

OpenStudy (highschoolmom2010):

22/7*0.5*0.5.4 what is that 22/7 part for

OpenStudy (nirmalnema):

it is replaced by pi

OpenStudy (highschoolmom2010):

im unsure how u got that

OpenStudy (nirmalnema):

it has a fixed value.... 22/7=3.14 something u can use that also

OpenStudy (highschoolmom2010):

so when u say the end value is pi V=3.14 ?

OpenStudy (highschoolmom2010):

im not sure how to write what the answer

OpenStudy (highschoolmom2010):

or do i need to leave it as pi yards ^3

OpenStudy (nirmalnema):

yup the volume= 3.14 or 22/7

OpenStudy (nirmalnema):

yes you can leave it like that also... because \[\pi * 0.5*0.5*4 = \pi * 1 =\] pi

OpenStudy (highschoolmom2010):

thanks @nirmalnema

OpenStudy (nirmalnema):

your welcome